OpenStudy (anonymous):

x+8/7 - x+3/6 (/ is fraction) Subtract and write in simplest form

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4fc1b012e4b0964abc836d0c use perenthesis to make it clearer

OpenStudy (anonymous):

it will be (8/7)+(3/6)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

so is anyone gonna help me or just tell me how to write it?

OpenStudy (anonymous):

it would be x-x=0 and 8/7+3/6=69/42

OpenStudy (anonymous):

and 69/42=23/14

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

tanx dear